OpenStudy (anonymous):

can someone help with the substitution rule for integrals? int:(3x+2)^2.4 dx

OpenStudy (jit4won):

put 3x+2 = t

OpenStudy (jit4won):

dx = dt/3

OpenStudy (jit4won):

I = int t^2.4 dt/3 = t^ 3.4 /10.2 + c

OpenStudy (anonymous):

what happens to the exponent? thats the only part i dont get

OpenStudy (jit4won):

Add 1

OpenStudy (jit4won):

2.4 + 1 = 3.4

OpenStudy (anonymous):

ok